# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 4.3. Chazy VI and $\mathcal{G}_{\mathrm{VI}}$
  • $q_{3}=[6:6:12]$, a saddle with eigenvalues $-5:6$.
At $q_{2}$, $B$ has a node, and $C$ has a smooth branch tangent to the branch of $B$ in the direction of the eigenvalue $1$; the curves $B$ and $C$ have a contact of order $6$ at $q_{2}$. The curve $C$ has a node at $q_{1}$, and $B$ passes also through $q_{3}$.
Upon resolving the foliation at $q_{2}$, we find a chain of five invariant rational curves of self-intersection $-2$, given by the four invariant components in the desingularization of $q_{2}$ plus the strict transform of $B$; see the left-hand side of Figure 7. Its contraction gives a singularity of type $A_{5}$ which, together with $p_{1}$ and $p_{2}$, gives the three singularities in the quotient model of $\mathcal{F}_{6}$.
@@
For $\Sigma=B^{-1}(1)$, we have a map $\Phi_{6}:\mathrm{Bl}_{3}(\mathbb{P}^{2})\to\Sigma$ [Gui12, p. 73], equivariant with respect to the actions of $\mathbb{Z}_{6}$ given the action of sixth roots of unity on $\Sigma$ via (7), and the action of $T_{6}$ on $\mathrm{Bl}_{3}(\mathbb{P}^{2})$ described in Section 3.3. It maps $\mathcal{E}_{6}$ to the foliation induced by $W$, and induces an identification of $\mathcal{G}_{\mathrm{VI}}$ with $\mathcal{F}_{6}$. From the explicit expressions for $\Phi_{6}$ and the inverse of $\Phi_{3}$ in [Gui12, p. 71], we may realize an explicit two-to-one map from $\mathbb{P}(1,2,3)$ to itself, mapping $\mathcal{G}_{\mathrm{IV}}$ to $\mathcal{G}_{\mathrm{VI}}$:
$[x:y:z]\mapsto[xz-2y^{2}:3x^{4}y^{2}-3x^{3}yz-9x^{2}y^{3}+x^{2}z^{2}+8xy^{2}z+4y^{4}-3yz^{2}:\\ 9x^{6}y^{3}-12x^{5}y^{2}z-45x^{4}y^{4}+6x^{4}yz^{2}+63x^{3}y^{3}z-x^{3}z^{3}+\\ +54x^{2}y^{5}-42x^{2}y^{2}z^{2}-48xy^{4}z+15xyz^{3}-16y^{6}+18y^{3}z^{2}-3z^{4}].$
It can be promoted to a rational map of $\mathbb{C}^{3}$ onto itself mapping the vector field corresponding to the Chazy IV equation to that of the Chazy VI one. This is the content of Theorem 8.
4.3.1. Birational equivalence with $\mathcal{F}_{6}$
Under the birational map (14), $\mathcal{G}_{\mathrm{VI}}$ is mapped to the degree two foliation of $\mathbb{P}^{2}$ given by (3). This establishes Theorem 3. That the foliation is the unique one of degree two on $\mathbb{P}^{2}$ tangent to both cubics follows, as before, from the fact that the tangency divisor of a pair of degree-two foliations on $\mathbb{P}^{2}$ has degree five.
The singularities in the complement of $x=0$ are those coming from $q_{1}$, $q_{2}$ and $q_{3}$, placed respectively at $(1:0:0)$, $(1:1:2)$ and $(18:3:1)$, with the same local descriptions as before. On the invariant line $x=0$, we have the singular point $(0:1:5)$, a saddle with eigenvalues $-1:2$, and $(0:0:1)$, a nilpotent singularity with multiplicity three. Its resolution is the composition of the resolution $\pi:S\to\mathbb{P}(1,2,3)$ with the map $j$ in (14). See the bottom-right of Figure 7.
@@ 5. Another plane model for Brunella’s foliation and its flop
In this section we will prove Theorem 9, that the degree two foliation $\mathcal{J}$ on $\mathbb{P}^{2}$ given by the form $\Omega$ in Eq. (10) is a planar model for Brunella’s very special foliation $\mathcal{F}_{3}$. In Section 5.2, we shall give another proof of the fact that the involutive Cremona map (11) represents the foliated flop.
The foliation $\mathcal{J}$ is tangent to the nodal cubic $C_{3}:xy^{2}-yz^{2}+zx^{2}-3xyz=0$, as well as to the lines of the coordinate triangle $\Delta_{3}:xyz=0$, each one of which is tangent to the cubic. Since a pair of degree-two foliations are either tangent along a curve of degree five or coincide, this is the only degree-two foliation tangent to this configuration. The singularities of $\mathcal{J}$ are:
  • $(0:0:1)$ , $(0:1:0)$ and $(1:0:0)$, non-degenerate, with eigenvalues $1:2$, and which are linearizable, for they have two tangent invariant curves through them: $C_{3}$ and one of the coordinate lines;
  • $(1:1:1)$, a reduced non-degenerate singularity with eigenvalues $-\omega:1$;
  • $(0:2:1)$, $(2:1:0)$ and $(1:0:2)$, which are reduced and non-degenerate, have eigenvalues $-3:2$, and are not on $C_{3}$.
A (minimal) reduction of singularities of $\mathcal{J}$ is given by six blow ups: at each of the vertices $(0:0:1)$, $(0:1:0)$, $(1:0:0)$, and along the infinitely near points along the directions of the local branches of the cubic $C_{3}$. This is schematically presented in Figure 8.
Figure 8. Reduction of singularities of $J$. Small gray curves are local separatrices.
In the blown-up projective plane in six points, the nodal curve has self-intersection $C_{3}^{2}=9-2-2-2=3$. The strict transforms of the lines of the triangle $\Delta_{3}$ and of the firstly introduced exceptional lines form three chains of two $(-2)$-curves in the blown-up plane (three $\mathcal{J}$-chains in the sense of [Bru15, Def. 8.1]), matching the combinatorics of the desingularization of a singularity of type $A_{2}$ each, as discussed in Section 2.2.
First proof of Theorem 9.
After blowing up, along $C_{3}$, there is just one reduced singularity, with eigenvalues $-\omega^{2}:1$. According to [Bru15] (Proposition 4.3), this characterizes the foliation $\mathcal{F}_{3}$ (up to birational equivalence). ∎
Second proof of Theorem 9.
@@
The birational equivalence of this proof is schematically described in Figure 9.
@@
Third proof of Theorem 9.
Let us give a proof in the spirit of the study on the Chazy equations carried out in [Gui12]. Consider the quadratic homogeneous vector field on $\mathbb{C}^{3}$
@@ 5.1. The de Jonquières symmetry
The Jacobian determinant of (11) is
@@
its fixed curve is the quartic
$x^{2}y^{2}+y^{2}z^{2}+x^{2}z^{2}-xy^{3}-yz^{3}-zx^{3}=0,$
of geometrical genus two, having a node at $(1:1:1)$. The involution $J_{4}$ preserves the lines through $(1:1:1)$, and is thus of de Jonquières type. The homaloidal system of $J_{4}$ is formed by the quartic plane curves with an ordinary triple point in $(1:1:1)$, and tangencies at $(0:0:1)$, $(0:1:0)$, and $(0:0:1)$. In the sense of [NTN20], this de Jonquières map is of type $(4;3;1^{2};1^{2};1^{2})$, of type 78.1 in Table 5.1, p. 102.
For the effect of $J_{4}$ on $\Omega$,
it follows that $J_{4}$ is indeed a birational automorphism of the foliation $\Omega=0$.
Figure 10 shows the elimination of indeterminations of the de Jonquières map: on top, the projective plane blown-up seven times is portrayed, and the two $(-1)$-curves which are interchanged under the flop are singled out (compare with Figure 4).
Figure 10. Diagram of elimination of indeterminations of $J_{4}$. The color of the arrow is the color of the exceptional line introduced by the blow-up.
In what follows, we will show that $J_{4}$ is the composition of three standard quadratic maps. This, and the proximity graph of Type 78.1, will show that its ordinary quadratic length (in the sense of [NTN20]) is three.
@@ 5.2. Factorization of the quartic de Jonquières involution
M. Noether established that every birational transformation of the complex projective plane may be factored as a composition of standard quadratic Cremona involutions and linear automorphisms. The varied and peculiar ways in which Cremona involutions may be composed to produce birational maps of small degrees bears witness to the complexity of this factorization; we refer the reader to recent works on the classification and factorization Cremona maps of degree three and four ([CD13], [CNTN22], [NTN20]) for a direct exposure to these.
We think that it is of interest to explicitly show that the quartic de Jonquières symmetry $J_{4}$ of $\mathcal{J}$ of Eq. (11) may be obtained as a composition of three standard quadratic Cremona maps; en passant, we shall give two models of $\mathcal{F}_{3}$ which are degree three foliations. In our planar model $\mathcal{J}$ for $\mathcal{F}_{3}$, a nodal plane cubic represents the link. So, in order to represent the foliated flop as a Cremona transformation of the plane, it is reasonable to try to obtain it as a composition of three quadratic Cremona maps which:
@@
This plan is carried out in what follows. Through it, we shall prove once again that the involutive Cremona map (11) represents the foliated flop. The choices in the changes of coordinates that follow were guided by Proposition 12.
5.2.1. The first quadratic map and its effect on the foliation
@@
Figure 11. Effect of the first quadratic Cremona map on the foliation; the indetermination points are in red and blue. The representation of the node is schematic; it is an acnode in the real plane.
@@
  • at $(0:0:1)$, $(0:1:0)$, $(-1:1:0)$: radial points (indicated by $1:1$ in Figure 11);
  • at $(0:1:-\omega)$, $(0:1:-\omega^{2})$, $(-1:0:1)$, $(1:-1:1)$, and $(-1:2:1)$: five non-degenerate singularities (the first two are the intersection $C_{2}\cap D_{1}$);
  • at $(-1:1:1)$: with eigenvalues $1:2$, linearizable;
@@
5.2.2. The second quadratic map and its effect on the foliation
Consider now the linear map $L_{2}(x:y:z):=(x-y+z:y-z:-z),$ which fixes $(1:0:0)$, and maps $(0:0:1)$ and $(0:1:0)$ to $(-1:1:1)$ and $(-1:1:0)$, respectively. The strict transform of $\mathcal{J}^{\prime}$ by the quadratic map $Q\circ L_{2}^{-1}$ is the degree three foliation $\mathcal{J}^{\prime\prime}$ given by the vanishing of
@@ 5.2.2. The second quadratic map and its effect on the foliation
Figure 12. Effect of the second quadratic Cremona map on the foliation; the indetermination points are in red and magenta.
@@
5.2.3. The third quadratic map and its effect on the foliation
@@
The line $C_{1}:x-y+z=0$ is contracted to the point $(1:0:0)$ by this third quadratic Cremona map and the strict transform of the conic $D_{2}:yz-xz+xy=0$ is the nodal cubic $D_{3}:xyz+xz^{2}+y^{2}z+y^{2}x=0$, having a node at $(1:0:0)$. See Figure 13.
Figure 13. Effect of the third Cremona map on the foliation. The indetermination points are in black and green.
@@
Finally, composing the three quadratic transformations $Q\circ L_{j}^{-1}$, $j=1,2,3$, and $L_{4}^{-1}$, we obtain (after extracting common factors) the map $J_{4}$ of Eq. (11), $J_{4}=L_{4}^{-1}\circ Q\circ L_{3}^{-1}\circ Q\circ L_{2}^{-1}\circ Q\circ L_{1}^{-1}$.
@@ 6. Quotients of linear foliations by standard quadratic Cremona involutions
The exceptionality of the automorphisms of linear foliations leading to the foliations of Brunella and Santos may be also brought to light through the study of the birational symmetries of linear foliations; we will study this in Section 7. There, we will also see that most linear foliations of the plane have only one non-linear birational automorphism (up to linear conjugation), the standard quadratic Cremona involution. The question of understanding the associated quotient foliations follows naturally. In this section, we will prove Theorem 12, giving explicit plane models for these foliations.
@@ 6.1. Cayley’s nodal cubic as the quotient under the standard quadratic involution
It has four singularities, which are nodal points (this is, they are of type $A_{1}$), at $(1:0:0:0)$, $(0:1:0:0)$, $(0:0:1:0)$ and $(0:0:0:1)$. It is the only cubic surface having four nodal points. It contains nine lines: six connecting a pair of nodes each, forming a tetrahedron; and three, coplanar, connecting pairs of points within the triple $(1:1:-1:-1)$, $(1:-1:1:-1)$, $(1:-1:-1:1)$, the lines $\xi+\eta=\zeta+\theta=0$, $\xi+\zeta=\eta+\theta=0$, and $\xi+\theta=\eta+\zeta=0$. The surface has no other lines. The group $S_{4}$ acts on $M_{3}$ by permuting the homogeneous coordinates, and naturally permutes all the above objects. See Figure 14.
@@
We will begin by showing that the quotient of $\mathbb{P}^{2}$ under the action of the standard quadratic Cremona involution $Q(X:Y:Z)=(YZ:ZX:XY)$ is Cayley’s nodal cubic surface (see also [Emc26, Section II.A]). Consider the $Q$-invariant rational map $\pi:\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{3}$,
@@
whose coordinates form a basis for the space of homogeneous cubic polynomials $P$ such that $P(YZ,ZX,XY)=XYZ\,P(X,Y,Z)$. It realizes the quotient of $\mathbb{P}^{2}$ by $Q$.
Let $\overline{\pi(\mathbb{P}^{2})}$ be the surface in $\mathbb{P}^{3}$ which is the rational image of $\mathbb{P}^{2}$ by $\pi$. It has the cubic equation
$\overline{\pi(\mathbb{P}^{2})}:\,\theta(\xi^{2}+\eta^{2}+\zeta^{2})-2\xi\eta\zeta-\theta^{3}=0.$
The four fixed points of $Q$ produce, via $\pi$, four singular points of type $A_{1}$ for $\overline{\pi(\mathbb{P}^{2})}$, placed at $(-1:-1:1:1)$, $(-1:1:-1:1)$, $(1:-1:-1:1)$ and $(1:1:1:1)$. The linear transformation
@@
establishes a linear isomorphism between $\overline{\pi(\mathbb{P}^{2})}$ and Cayley’s nodal cubic (29), and maps the above singular points to the nodes on Cayley’s surface.
The map $L\circ\pi$ realizes the quotient of $\mathbb{P}^{2}$ under the action of the standard quadratic Cremona involution as Cayley’s nodal cubic surface.
6.2. A birational plane model
Let us take a step further, and pass from Cayley’s surface to a new copy of the projective plane. The strict transform of Cayley’s cubic surface $M_{3}$ by the cubic involutive Cremona map of $\mathbb{P}^{3}$
@@ 6.2. A birational plane model
6.3. The quotient foliations

### v2

@@ 4.3.3. Description of $\mathcal{G}_{\mathrm{VI}}$
  • $q_{3}=[6:6:12]$, a saddle with eigenvalues 5:6-5:6.
At $q_{2}$, $B$ has a node, and $C$ has a smooth branch tangent to one of the branches of $B$ at it. This point is thus a linearizable singularity of the foliation, and the curves have a total contact of contact of order six at it. The curve $C$ has a node at $q_{1}$, and $B$ passes also through $q_{3}$.
Upon resolving the foliation at $q_{2}$, we find a chain of five invariant rational curves of self-intersection 2-2, given by the four invariant components in the desingularization of $q_{2}$, plus the strict transform of $B$; see the left-hand side of Figure 7. Its contraction gives a singularity of type $A_{5}$ which, together with $p_{1}$ and $p_{2}$, gives the three singularities in the quotient model of $\mathcal{F}_{6}$.
@@
4.3.4. Birational equivalence of $\mathcal{F}_{6}$ and $\mathcal{H}_{6}$
Under the birational map (14), $\mathcal{G}_{\mathrm{VI}}$ is mapped to the degree two foliation of $\mathbb{P}^{2}$ given by (3). The invariant curve $x=0$ for $\mathcal{H}_{6}$ is produced by the map (14); its other invariant curves are the strict transforms of $B$ and $C$. That the foliation is the unique one of degree two on $\mathbb{P}^{2}$ tangent to both cubics follows, as before, from the fact that the tangency divisor of a pair of degree-two foliations on $\mathbb{P}^{2}$ has degree five. This establishes Theorem 3.
The singularities of $\mathcal{H}_{3}$ in the complement of $x=0$ are those coming from the singularities $q_{1}$, $q_{2}$ and $q_{3}$ of $\mathcal{G}_{\mathrm{VI}}$; they are placed, respectively, at $(1:0:0)$, $(1:1:2)$ and $(18:3:1)$, and have the same local descriptions. On the invariant line $x=0$, we have the singular point $(0:1:5)$, a saddle with eigenvalues 1:2-1:2, and $(0:0:1)$, a nilpotent singularity with multiplicity three. Its resolution is the composition of the resolution $\varpi:S\to\mathbb{P}(1,2,3)$ with the map $j$ in (14). See the bottom-right of Figure 7.
@@ 5. Another plane model for Brunella’s foliation and its flop
In this section we will study the foliation $\mathcal{J}$ on $\mathbb{P}^{2}$ given by the form $\Omega$ of Eq. (10). Our aim is to establish Theorem 9, in particular, that it is a planar model for Brunella’s very special foliation $\mathcal{F}_{3}$. We describe the foliation in Section 5.1, and establish the aforementioned birational equivalence in Section 5.2. In Section 5.3, we will see that (11) is an involutive birational automorphism of $\mathcal{J}$, and that it corresponds to Brunella’s foliated flop. In Section 5.4, we shall give another proof of this last fact, along with a detailed factorization of the map (11) into quadratic Cremona maps.
5.1. Description of $\mathcal{J}$
The foliation is tangent to the nodal cubic $C_{3}:xy^{2}-yz^{2}+zx^{2}-3xyz=0$, as well as to the lines of the coordinate triangle $\Delta_{3}:xyz=0$, each one of which is tangent to the cubic. Since a pair of degree-two foliations are either tangent along a curve of degree five or coincide, this is the only degree-two foliation tangent to this configuration. The singularities of $\mathcal{J}$ are:
  • $(0:0:1)$ , $(0:1:0)$ and $(1:0:0)$, non-degenerate, with eigenvalues $1:2$, and which are linearizable, for they have two tangent invariant curves through them: $C_{3}$ and the coordinate line tangent to it at this point;
  • $(1:1:1)$, a reduced non-degenerate singularity with eigenvalues $\omega:1$;
  • $(0:2:1)$, $(2:1:0)$ and $(1:0:2)$, which are reduced and non-degenerate, have eigenvalues 3:2-3:2, and are not on $C_{3}$.
A minimal reduction of singularities of $\mathcal{J}$ is given by six blow ups: at each of the vertices $(0:0:1)$, $(0:1:0)$, $(1:0:0)$, and along the infinitely near points along the directions of the local branches of $C_{3}$. This is schematically presented in Figure 8.
Figure 8. Reduction of singularities of $\mathcal{J}$. Small gray curves are local separatrices.
In the blown-up projective plane in six points, the nodal curve, strict transform of $C_{3}$, has self-intersection $C_{3}^{2}=9-2-2-2=3$. The strict transforms of the lines of the triangle $\Delta_{3}$ and of the firstly introduced exceptional lines form three chains of two (2)(-2)-curves in the blown-up plane (three $\mathcal{J}$-chains in the sense of [Bru15, Def. 8.1]), matching the combinatorics of the desingularization of a singularity of type $A_{2}$ each, as discussed in Section 2.2.
5.2. The birational equivalence with $\mathcal{F}_{3}$
Let us establish the main fact of Theorem 9, that the degree two foliation $\mathcal{J}$ on $\mathbb{P}^{2}$ given by the form $\Omega$ in Eq. (10) is a planar model for Brunella’s very special foliation $\mathcal{F}_{3}$. We will give four proofs of it:
First proof.
After reduction of the singularities of $\mathcal{J}$, along the strict transform of $C_{3}$, there is just one reduced singularity, with eigenvalues $\omega^{2}:1$. According to [Bru15, Prop. 4.3], this characterizes the foliation $\mathcal{F}_{3}$ (up to birational equivalence). ∎
Second proof.
@@ 5.2. The birational equivalence with $\mathcal{F}_{3}$
The birational equivalence of this proof is schematically presented in Figure 9.
@@
Third proof.
Let us now give a proof in the spirit of the study on the Chazy equations carried out in [Gui12]. Consider the quadratic homogeneous vector field on $\mathbb{C}^{3}$
@@ 5.3. The de Jonquières symmetry
Let us now discuss the fact that the transformation $J_{4}$ of Eq. (11) is a birational involution of preserving $\mathcal{J}$, that corresponds to Brunella’s foliated flop. We have already calculated the flop in Pereira’s model (1), and through the explicit map of the second proof in the previous subsection, we may establish this fact. It may also be calculated from the automorphism $Q$ of Eq. (17), via our third proof. It can also be established by a direct calculation that $J_{4}$ is an automorphism of $\mathcal{J}$: for the effect of $J_{4}$ on $\Omega$,
and it follows that $J_{4}$ is indeed a birational automorphism of the foliation defined by $\Omega=0$. (The first item of Theorem 5 implies that it corresponds to Brunella’s foliated flop, up to a cyclic permutation of the coordinates.) Together with the previous results, this establishes Theorem 9.
We can give further information on $J_{4}$. Its Jacobian is
@@
its fixed curve is the quartic $x^{2}y^{2}+y^{2}z^{2}+x^{2}z^{2}-xy^{3}-yz^{3}-zx^{3}=0$, of geometrical genus two, having a node at $(1:1:1)$. The involution $J_{4}$ preserves the lines through $(1:1:1)$, and is thus of de Jonquières type. The homaloidal system of $J_{4}$ is formed by quartic plane curves with an ordinary triple point in $(1:1:1)$, and tangencies at $(0:0:1)$, $(0:1:0)$, and $(0:0:1)$. In the sense of [NTN20], this de Jonquières map is of type $(4;3;1^{2};1^{2};1^{2})$, of type 78.1 in Table 5.1, p. 102.
Figure 10 shows the elimination of indeterminacies of the de Jonquières map: on top, the projective plane blown-up seven times is portrayed, and the two (1)(-1)-curves which are interchanged under the flop are singled out (compare with Figure 4).
Figure 10. Diagram of elimination of indeterminacies of $J_{4}$. The color of the arrow is the color of the exceptional line introduced by the blow-up.
@@ 5.4. Factorization of the quartic de Jonquières involution
M. Noether established that every birational transformation of the complex projective plane may be factorized as a composition of standard quadratic Cremona involutions and linear automorphisms. A natural measure of the complexity of a birational map is thus the minimum number of standard quadratic maps appearing in such a factorization, called its ordinary quadratic length. It follows from the previously discussed facts that the ordinary quadratic length of $J_{4}$ is three, since $J_{4}$ is of type 78.1 in Table 5.1, p. 102 in [NTN20]. Our aim here is to present an explicit factorization of $J_{4}$ into standard quadratic transformations and linear ones, in a way realizing its ordinary quadratic length.
The varied and peculiar ways in which Cremona involutions may be composed to produce birational maps of small degrees bear witness to the complexity of Noether’s factorization; we refer the reader to recent works on the classification and factorization Cremona maps of degree three and four ([CD13], [CNTN22], [NTN20]) for a direct exposure to these.
The quartic de Jonquières symmetry $J_{4}$ Eq. (11) was not obtained by trial and error, but deliberately built as a composition of standard quadratic Cremona maps and linear automorphisms following some guidelines. Since in our planar model $\mathcal{J}$ for $\mathcal{F}_{3}$ a nodal plane cubic represents the link, then, in order to represent the foliated flop as a Cremona transformation of the plane, it seemed plausible to obtain $J_{4}$ as a composition of three quadratic Cremona maps, which:
@@
By choosing changes of coordinates guided by Proposition 12, we succeeded in doing this. The results are presented in what follows. As a consequence, we will establish once again that the involutive Cremona map (11) represents the foliated flop.
The first quadratic map and its effect on the foliation
@@
Figure 11. Effect of the first quadratic Cremona map on the foliation; the indeterminacy points are in red and blue. The representation of the node is schematic; it is an acnode in the real plane.
@@
  • at $(0:0:1)$, $(0:1:0)$, (1:1:0)(-1:1:0): radial points (indicated by $1:1$ in Figure 11);
  • at $(0:1:-\omega)$, $(0:1:-\omega^{2})$, (1:0:1)(-1:0:1), (1:1:1)(1:-1:1), and (1:2:1)(-1:2:1): five non-degenerate singularities (the first two are the intersection $C_{2}\cap D_{1}$);
  • at (1:1:1)(-1:1:1): with eigenvalues $1:2$, linearizable;
@@
The second quadratic map and its effect on the foliation
Consider now the linear map $L_{2}(x:y:z):=(x-y+z:y-z:-z),$ which fixes $(1:0:0)$, and maps $(0:0:1)$ and $(0:1:0)$ to (1:1:1)(-1:1:1) and (1:1:0)(-1:1:0), respectively. The strict transform of $\mathcal{J}^{\prime}$ by the quadratic map $Q\circ L_{2}^{-1}$ is the degree three foliation $\mathcal{J}^{\prime\prime}$ given by the vanishing of
@@
Figure 12. Effect of the second quadratic Cremona map on the foliation; the indeterminacy points are in red and magenta.
@@
The third quadratic map and its effect on the foliation
@@
The line $C_{1}:x-y+z=0$ is contracted to the point $(1:0:0)$ by this third quadratic Cremona map, and the strict transform of the conic $D_{2}:yz-xz+xy=0$ is the nodal cubic $D_{3}:xyz+xz^{2}+y^{2}z+y^{2}x=0$, whose node is at $(1:0:0)$. See Figure 13.
Figure 13. Effect of the third Cremona map on the foliation. The indeterminacy points are in black and green.
@@
Finally, composing the three quadratic transformations $Q\circ L_{j}^{-1}$, $j=1,2,3$, and $L_{4}^{-1}$, we obtain (after extracting common factors) the map $J_{4}$ of Eq. (11),
$J_{4}=L_{4}^{-1}\circ Q\circ L_{3}^{-1}\circ Q\circ L_{2}^{-1}\circ Q\circ L_{1}^{-1}.$
We have thus shown that $J_{4}$ can be factored as the composition of linear automorphisms and three ordinary quadratic Cremona transformations.
@@ 6. Quotients of linear foliations by standard quadratic Cremona involutions
The exceptionality of the automorphisms of linear foliations leading to the foliations of Brunella and Santos may be also brought to light through the study of the birational symmetries of linear foliations; we will study this in Section 7. There, we will also see that most linear foliations of the plane have only one non-linear birational automorphism (up to linear conjugation), the standard quadratic Cremona involution. The question of understanding the associated quotient foliations follows naturally. In this section, we will prove Theorem 12, giving explicit plane models for these quotient foliations.
@@ 6.1. Cayley’s nodal cubic as the quotient under the standard quadratic involution
It has four singularities, which are nodal points (this is, they are of type $A_{1}$), at $(1:0:0:0)$, $(0:1:0:0)$, $(0:0:1:0)$ and $(0:0:0:1)$. It is the only cubic surface having four nodal points. It contains nine lines: six connecting a pair of nodes each, forming a tetrahedron; and three, coplanar, connecting pairs of points within the triple (1:1:1:1)(1:1:-1:-1), (1:1:1:1)(1:-1:1:-1), (1:1:1:1)(1:-1:-1:1), the lines $\xi+\eta=\zeta+\theta=0$, $\xi+\zeta=\eta+\theta=0$, and $\xi+\theta=\eta+\zeta=0$. The surface has no other lines. The group $S_{4}$ acts on $M_{3}$ by permuting the homogeneous coordinates, and naturally permutes all of the above objects. See Figure 14.
@@
Our starting point will be the following fact. It has probably been known for a very long time, but the earliest reference we found was that of Emch [Emc26, Section II.A].
Proposition 18.
The quotient of $\mathbb{P}^{2}$ under the action of the standard quadratic Cremona involution $Q(X:Y:Z)=(YZ:ZX:XY)$ is Cayley’s nodal cubic surface.
Proof.
Consider the $Q$-invariant rational map $\pi:\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{3}$,
@@
whose coordinates form a basis for the space of homogeneous cubic polynomials $P$ such that P(YZ,ZX,XY)=XYZP(X,Y,Z)P(YZ,ZX,XY)=XYZ\,P(X,Y,Z). The rational image of the plane by $\pi$ is a (singular) complex surface $\pi(\mathbb{P}^{2})$ of $\mathbb{P}^{3}$ and $\pi$ is a quotient map for the action of $Q$. The surface $\pi(\mathbb{P}^{2})$ is given by the cubic equation
$\theta(\xi^{2}+\eta^{2}+\zeta^{2})-2\xi\eta\zeta-\theta^{3}=0.$
The four fixed points of $Q$ produce, via $\pi$, four singular points of type $A_{1}$ for $\pi(\mathbb{P}^{2})$, placed at (1:1:1:1)(-1:-1:1:1), (1:1:1:1)(-1:1:-1:1), (1:1:1:1)(1:-1:-1:1) and $(1:1:1:1)$. The linear transformation
@@
establishes a linear isomorphism between $\pi(\mathbb{P}^{2})$ and Cayley’s nodal cubic (31), and maps the above singular points to the nodes on Cayley’s surface. In this way, $L\circ\pi$ realizes the quotient of $\mathbb{P}^{2}$ under the action of the standard quadratic Cremona involution as Cayley’s nodal cubic surface. ∎
The finite quotients of the plane are rational surfaces, that is, birationally equivalent to the plane. Let us show an explicit equivalence for Cayley’s cubic. The strict transform of Cayley’s cubic surface $M_{3}$ by the cubic involutive Cremona map of $\mathbb{P}^{3}$
@@
6.2. The quotients of degree-one foliations
After these preliminaries and Proposition 19, we are ready to prove the sought result.
Proof of Theorem 10.